Some Bugs in the Travel Planning Corpus

Manuel Kauers

November 23, 2001

1 Assumptions

I have closely considered the file cmu-german.db as offered by Stefan Vogel, and I met the following curiosities which in my opinion are likely to be spelling bugs.

Due to a lack of a rigorous IL specification, I have assumed the following grammar (in EBNF):

⟨domain-action⟩ → ⟨speaker⟩':'⟨speech-act⟩⟨concept-list⟩⟨argument-list⟩
⟨speaker⟩ → ⟨identifier⟩
⟨speach-act⟩ → ⟨identifier⟩
⟨concept-list⟩ → ( '+'⟨concept⟩ )^{∗}
⟨concept⟩ → ⟨identifier⟩
⟨argument-list⟩ → '' | '('⟨argument⟩ ( ','⟨argument⟩ )^{∗} ')'
⟨argument⟩ → ⟨arg-name⟩'='⟨arg-value⟩
⟨arg-name⟩ → ⟨identifier⟩
⟨arg-value⟩ → ⟨atomic⟩ | ⟨multiple⟩ | ⟨coordination⟩ | ⟨special⟩
⟨atomic⟩ → ⟨identifier⟩
⟨multiple⟩ → '('⟨arg-or-at⟩ ( ','⟨arg-or-at⟩ )^{∗} ')'
⟨coordination⟩ → '('⟨arg-or-at⟩ ( '&'⟨arg-or-at⟩ )^{∗} ')'
⟨special⟩ → '['⟨arg-or-at⟩ ( ','⟨arg-or-at⟩ )^{∗} ']'
⟨arg-or-at⟩ → ⟨atomic⟩ | ⟨argument⟩
⟨identifier⟩ → ⟨any-sequence-of-characters-or-numbers-or-hyphens⟩

Here, ⟨any-sequence-of-characters-or-numbers-or-hyphes⟩ expands to an arbitrary sequence of characters (such as 'a', 'b', 'c', …), numbers (such as '0', '1', '2', …), and hyphens (such as '-', ' '). White spaces where allowed before and after ':', '+', '(', ')', ',', '&', '[', and ']', but not within identifiers.

In 6.10.1, 22.50.1, and 28.9.1, I observed parentheses around arguments themselves, i.e. something like

a : b+c+d(e = f, \underline{(} g = h \underline{)} ).

I suppose that this was not meant, because I cannot imagine of a semantics for that.

2 Syntactic problems

location there is …  there should be … / remarks
10.15.1 who = i e-time  who = i, e-time
11.2.3 ++concept +concept
12.40.1 c::  c:
17.35.1 , , location  , location
22.43.1 c:(  c:
23.14.1 tour-spec-(  tour-spec = (
25.24.1 [spouse  [(spouse
25.24.1 )])  )]))
25.69.3 = * (  = (
44.52.1 focalizer-additive  focalizer = additive
44.85.1 feature full board  feature = full board
44.76.2 c:(  c:
44.107.1 eventuality = ,  eventuality = ??,
44.114.2 a:no-tag + a:no-tag / parser knock out ;-)
44.130.4 hotel identifiability hotel, identifiability
44.191.3 spec-(  spec = (
44.195.2 spec-(  spec = (
50.34.3 spec-(  spec = (
50.51.1 rhetorical-condition rhetorical = condition
50.89.2 yes). time  yes), time
50.103.5 yes) feature  yes), feature
55.72.6 location = ,  location = ??,
55.106.1 info-topic-(  info-topic = (

3 Semantic problems

I have also discovered some strange identifiers which are probably simply misspelled.

location there is …  there should be … / remarks
10.52.1  additve  additive
17.38.2  identifiabiility  identifiability
20.10.3  adamello brenta park adamello-brenta park
44.8.2  identifability  identifiability
44.16.2  identiability  identifiability
44.26.7  identiability  identifiability
44.178.10 identiability  identifiability
44.167.1 identifiablity  identifiability

50.11.1  desre  desire
50.47.1  accomdation-class  accomodation-class
50.47.1  reservartion  reservation
50.51.1  identifiablity  identifiability
50.80.2  halfboard  half board
50.83.3  idenfitifiability  identifiability

Greetings from Karlsruhe, Manuel